OpenStudy (anonymous):

Create two sets of three triangle lengths: one set that does create a right triangle and another that does not. Show and explain why the first set does create a right triangle and why the second set does not.

OpenStudy (anonymous):

first set is h=50 ,b=30; p=40; h=sqrt(30^2+40^2) =50 because this follow the pythagorian law. which hold only for right triangle. other set of trianle h=3; b=2;p=5 h=sqrt(2^2+5^2) is not equal to 3